# Unsupervised Meta-Learning via In-Context Learning

### v2

@@ Abstract
Unsupervised meta-learning aims to learn feature representations from unsupervised datasets that can transfer to downstream tasks with limited labeled data. In this paper, we propose a novel approach to unsupervised meta-learning that leverages the generalization abilities of in-context learning observed in transformer architectures. Our method reframes meta-learning as a sequence modeling problem, enabling the transformer encoder to learn task context from support images and utilize it to predict query images. At the core of our approach lies the creation of diverse tasks generated using a combination of data augmentations and a mixing strategy that challenges the model during training while fostering generalization to unseen tasks at test time. Experimental results on benchmark datasets showcase the superiority of our approach over existing unsupervised meta-learning baselines, establishing it as the new state-of-the-art in the field. Remarkably, our method achieves competitive results with supervised and self-supervised approaches, underscoring the efficacy of the model in leveraging generalization over memorization.
**footnotetext: Equal contribution
@@ 1 Introduction
Meta-learning, or learning-to-learn, empowers learning systems with the ability to accumulate knowledge from multiple tasks, enabling rapid adaptation and generalization to new tasks [50, 47]. Traditional meta-learning approaches typically rely on labeled data to construct tasks during meta-training. However, collecting large labeled datasets in real-world applications is challenging and often impractical. Addressing this issue, unsupervised meta-learning (UML) approaches aim to leverage unlabeled data to learn transferable feature representations, enabling adaptation to new tasks with limited labeled data [50].
@@
Various approaches have been proposed to address the UML problem [19, 20, 21, 24, 25, 26]. However, UML still faces several challenges. Existing UML methods often rely on simple data augmentations to construct the training tasks, while following the standard meta-learning task sampling pipeline for the evaluation phase. This results in a significant difference between training and testing tasks, constraining the model’s ability to generalize to unseen tasks during inference and requiring additional fine-tuning on the test domain. Furthermore, existing UML approaches typically assume that the training and test datasets belong to the same domain. In our framework, we loosen this assumption resulting in a more challenging setting that necessitates a better model generalization compared to usual meta-learning applications. We refer to this as the cross-domain scenario.
@@
In this paper, we propose a novel approach to UML that addresses these challenges by leveraging in-context learning within a transformer architecture [11, 32]. In-context learning allows the model to use the context provided by a sequence of input-output pairs to make predictions on new input data. Inspired by recent advancements in large language models (LLMs) [54, 4, 30], we formulate meta-learning as a sequence modeling problem, where a task is seen as a sequence of support images and an unknown query image. The support set is treated as the context utilized by the model to predict the class of the query image. We call our approach CAMeLU, which stands for Context-Aware Meta-Learning in Unsupervised scenarios. Central to our approach is a novel task creation mechanism that enables the generation of a large number of different tasks from an unlabeled dataset. Drawing inspiration from the natural decision process of learning by analogy [56], we construct tasks that closely resemble the structure of those encountered during inference. Specifically, we use a combination of different data augmentation techniques based on basic image manipulations [41] for generating the samples in the support set. Conversely, a strategy similar to mixup [60] is employed to generate query images by combining a support element, after applying a distinct augmentation function, and an image randomly sampled from the training dataset. This process ensures that the query contains sufficient information from the support image to be classified as the latter, while introducing diversity by blending them. Consequently, query images appear distinct from their corresponding support images while still belonging to the same class, better mimicking the tasks seen at test time and hence enhancing generalization. Following task creation, support and query images are encoded using a fixed pre-trained feature extractor. The resulting latent representations are aggregated into a sequence and passed as input to a transformer encoder along with their label encodings. The transformer encoder learns to extract contextual information from support images and predict the query image in a single pass, eliminating the need for the fine-tuning step during inference. An overview of our approach is visualized in Figure 1.
@@
Figure 1: Visualization of CAMeLU (with 3-way 5-shot tasks). The left side illustrates the task creation mechanism, where $N$ samples are drawn from an unlabeled dataset $\mathcal{D}_{train}$. Each sample $x_{n}$ is augmented $K$ times to obtain $x_{n,k}^{(sp)}$. A strategy inspired by mixup [60] is utilized for generating the query set by using an augmented version of $x_{n}$, i.e., $\tilde{x}_{n,j}$. The same pseudo-label $n\in[1,N]$ is assigned to all data generated from the sample $x_{n}$. On the right side, the so-created task is fed into the transformer encoder for predicting the query input. Inspired by CAML [12], the transformer encoder processes demonstrations created by concatenating features from a fixed pre-trained feature extractor and a learned class encoder. The symbol * denotes the unknown query label that the transformer encoder aims to predict.
@@ 2 Related work
Meta-learning is a well-studied field in the machine learning community due to its ability to enable models to quickly adapt to tasks with limited labeled data. Pioneering work in the field [14, 42, 51, 33, 43] considers the scenarios where a large labeled dataset is available for meta-training, a challenging requirement in real-world applications. UML addresses this challenge by extracting meaningful information from unsupervised data that can be transferred to downstream tasks with limited labeled data. Different techniques have been explored in the literature to construct diverse tasks. CACTUs [19] applies clustering in the embedding space and assigns the same pseudo-label to all images in the same cluster. Other methods focus on generating synthetic samples, either using data augmentations, as in UMTRA [21], or leveraging interpolation in the latent space of a generative model [22]. Differently, Meta-GMVAE [26] and Meta-SVEBM [24] use variational autoencoders and memory-based models for pseudo-label generation. Recent methodologies have also incorporated SSL techniques [10] into UML methods. In particular, Set-SimCLR [25] builds on top of the SimCLR [7] approach and reframes meta-learning as a set-level problem, while PsCo [20], inspired by MoCo [16], utilizes a momentum encoder and a queue of previous samples to improve pseudo-labeling and construct diverse tasks for UML applications. Similarly, BECLR [34] introduces an approach for unsupervised few-shot learning by proposing a constrastive representation learning framework, instead of meta-learning.
@@
Several UML approaches rely on data augmentation to construct the training tasks [21, 25, 20]. However, traditional transformations such as rotation, translation, cropping, resizing, and flipping [41] might generate images that are too similar to the original ones, ending up in tasks with low in-class variability between the support and query images. This creates a problem when the model needs to generalize to test tasks, where the query data are different instances than the support ones, not only augmented versions of them. In this paper, we addressed this limitation by generating query images using a strategy inspired by mixup [60] to enhance model generalization. Similarly to mixup, which performs linear interpolation of the feature vectors at the pixel level, other strategies based on mixing images comprise CutMix [59], PatchMix [29], and Manifold Mixup [49].
@@
In-context learning refers to the ability to perform a new task via inference alone by conditioning on a few input-output pairs and making predictions for new inputs [11]. While typical of LLMs [9, 37, 44], this ability has also been explored in different fields, such as in-painting [1, 61], image segmentation [5], and notably meta-learning [23, 13, 12, 32]. In particular, GPICL [23] shows that transformers can be meta-trained to act as general-purpose in-context learners. Similarly, CAML [12] proposes a universal meta-learner by rephrasing the meta-learning framework as a sequence modeling problem and employing a transformer model for few-shot classification. However, this approach is designed for supervised meta-learning, thus requiring labels both during the training and inference phases. Inspired by these, we propose a new approach that integrates the in-context learning ability of transformers with a novel task creation mechanism to solve the UML problem.
@@ 3 Proposed approach
Our proposed approach, Context-Aware Meta-Learning in Unsupervised scenarios (CAMeLU), leverages the in-context learning ability of transformers to address the challenges of UML. These challenges include the need to construct meaningful tasks from unlabeled data and the requirement for models to generalize effectively to new tasks during inference. CAMeLU consists of two phases that are intertwined during the model training. Initially, tasks are automatically constructed from an unlabeled dataset utilizing a combination of two strategies. Subsequently, we reformulate the meta-learning framework as a sequence modeling problem, aiming to harness the in-context learning capability of a transformer. This enables the model to extract context from the support samples and predict the unknown query samples without requiring any fine-tuning during the inference phase. We delve into the two phases in Sect. 3.1 and Sect. 3.2, respectively.
@@ 3.1 Task creation
For the support set, we randomly sample $N$ images from $\mathcal{D}_{train}$ under the assumption that they belong to distinct categories, as shown in Figure 1. This assumption is reasonable when $N<<C$, where $C$ denotes the total number of classes in $\mathcal{D}_{train}$, which is satisfied using a large training dataset. If we assume that all samples are equally distributed among the classes, i.e., $m$ samples per class, the probability that two or more samples are in the same class is equal to
@@
This task-creation mechanism can be seen as a task augmentation strategy [58, 38] that allows the generation of a large number (almost infinite) of diverse tasks. This is particularly useful for meta-learning and in-context learning applications where the model needs to acquire knowledge from a multitude of tasks to generalize to unseen tasks sampled from different domains.
@@
While the strategy used for generating the query images draws inspiration from the mixup strategy proposed in [60], there are some substantial differences. The aim of mixup is to develop a new data augmentation strategy to expand the number of training examples and diversify the data distribution used for training, thereby enhancing the robustness and generalization of neural networks. In CAMeLU, the primary objective of merging images is to encourage the model to learn even in scenarios where only a fraction of the class context is present in the image. In CAMeLU, $\lambda$ is sampled from a uniform distribution (obtained with a $Beta$ distribution with $\alpha=1,\beta=1$) in $(0,0.5)$, guaranteeing that the amount of information from $z_{j}$ that is embedded into $x_{j}^{(qr)}$ is less than 50%, thus ensuring that the assigned label is consistent with the class of the support images generated from $x_{n}$. Indeed, we assign the same label $n$ to $x_{j}^{(qr)}$, forcing the network to learn to retrieve information in $x_{j}^{(qr)}$ that is related to the category of $x_{n}$. Contrarily, mixup creates new examples by interpolating both images and labels at the scope of limiting memorization over the training distribution.
@@ 3.2 In-context learning method
Following task creation, we rephrase the meta-learning framework as a sequence modeling problem. Inspired by recent developments in LLMs [15, 27, 9, 37, 44], we treat each task as a prompt, where the support embedding, together with the learned projected labels, form the demonstration context, whereas the query represents the classification problem that the network is required to solve. A model is said to in-context learn a task if it can approximate $y_{j}^{(qr)}$ for a new query input $x_{j}^{(qr)}$ by conditioning on a sequence $S_{i,j}$ containing in-context (support) examples and one query input defined as follows:
@@
To achieve this, we design a model comprising three components: (1) a feature extractor $f_{\psi}$, (2) a class encoder $g_{\phi}$, and (3) a transformer encoder with a linear projection layer on top, i.e., $M_{\theta}$. The feature extractor aims to map support and query samples into a latent space where images with similar characteristics and semantic meaning are assigned similar representations. In Appendix A.4 we explore various feature extractors for this purpose, including those pre-trained via a supervised approach or leveraging an SSL technique. The resulting representations are then concatenated with a class embedding. The class embeddings for the support representations are generated by encoding the corresponding classes using the class encoder $g_{\phi}$. However, as the classes of the queries are unknown, a randomly initialized learnable vector is appended to each query representation. The so-combined embeddings are then organized into sequences $S_{i,j}=\left((f_{\psi}(x_{1}^{(sp)}),g_{\phi}(y_{1}^{(sp)})),\ldots,(f_{\psi}(x_{NK}^{(sp)}),g_{\phi}(y_{NK}^{(sp)})),f_{\psi}(x_{j}^{(qr)})\right),$ $j=1,\ldots,Q,$ resembling the one in Equation 2. These sequences are fed into the transformer encoder, and only the transformer output corresponding to the query sample is selected and passed through a projection layer to predict the query label. This process iterates for all queries in the task, and the aggregated loss is employed for model training. In particular, the training process can be formulated as an optimization process where the objective is as follows:
@@
During evaluation, when a new task is presented, the available examples in $D_{new}^{(sp)}$ are utilized as contextual information to guide the classfication of the query samples, without requiring any fine-tuning or adaptation steps.

### v3

@@ Abstract
Unsupervised meta-learning aims to learn feature representations from unsupervised datasets that can transfer to downstream tasks with limited labeled data. In this paper, we propose a novel approach to unsupervised meta-learning that leverages the generalization abilities of in-context learning observed in transformer architectures. Our method reframes meta-learning as a sequence modeling problem, enabling the transformer encoder to learn task context from support images and utilize it to predict query images. At the core of our approach lies the creation of diverse tasks generated using a combination of data augmentations and a mixing strategy that challenges the model during training while fostering generalization to unseen tasks at test time. Experimental results on benchmark datasets showcase the superiority of our approach over existing unsupervised meta-learning baselines, establishing it as the new state-of-the-art. Remarkably, our method achieves competitive results with supervised and self-supervised approaches, underscoring its efficacy in leveraging generalization over memorization.
@@ 1 Introduction
Meta-learning, or learning-to-learn, enables models to accumulate knowledge from multiple tasks, allowing rapid adaptation and generalization to new tasks (Vettoruzzo et al., 2024; Vanschoren, 2019). Traditional meta-learning approaches typically rely on labeled data to construct tasks during meta-training. However, collecting large labeled datasets in real-world applications is challenging and often impractical. Unsupervised meta-learning (UML) methods address this issue by leveraging unlabeled data to learn transferable feature representations, enabling adaptation to new tasks with limited labeled data (Vettoruzzo et al., 2024).
@@
Various approaches have been proposed to address the UML problem (Hsu et al., 2018; Jang et al., 2022; Khodadadeh et al., 2019; Kong et al., 2021; Lee et al., 2022; 2020). However, UML still faces several challenges. Existing UML methods often rely on simple data augmentations to construct the training tasks, while following the standard meta-learning task sampling pipeline for evaluation. This results in a significant difference between training and testing tasks, limiting generalization and often requiring fine-tuning on the test domain. Furthermore, existing UML approaches typically assume that the training and test datasets belong to the same domain. In our framework, we loosen this assumption resulting in a more challenging setting that necessitates a better model generalization compared to usual meta-learning applications. We refer to this as the cross-domain scenario.
@@
In this paper, we propose a novel approach to UML that addresses these challenges by leveraging in-context learning within a transformer architecture (Dong et al., 2022; Min et al., 2022). In-context learning allows the model to use the context provided by a sequence of input-output pairs to make predictions on new input data. Inspired by recent advancements in large language models (LLMs) (Wei et al., 2022; Brown et al., 2020; Liu et al., 2022), we formulate meta-learning as a sequence modeling problem, where a task is seen as a non-causal sequence of support images and an unknown query image. The support set is treated as the context utilized by the model to predict the class of the query image. We call our approach CAMeLU, which stands for Context-Aware Meta-Learning in Unsupervised scenarios. Central to our approach is a novel task creation mechanism that enables the generation of a large number of different tasks from an unlabeled dataset. Drawing inspiration from the natural decision process of learning by analogy (Winston, 1980), we construct tasks that closely resemble the structure of those encountered during inference. Specifically, we use a combination of different data augmentation techniques based on basic image manipulations (Shorten & Khoshgoftaar, 2019) for generating the samples in the support set. Conversely, a strategy similar to mixup (Zhang et al., 2018) is employed to generate query images by combining a support element, after applying a distinct augmentation function, and an image randomly sampled from the training dataset. This process ensures that the query contains sufficient information from the support image to be classified as the latter, while introducing diversity by blending them. Consequently, query images appear distinct from their corresponding support images while still belonging to the same class, better mimicking the tasks seen at test time and hence enhancing generalization. Following task creation, support and query images are encoded using a fixed pre-trained feature extractor. The resulting latent representations are aggregated into a sequence and passed as input to a transformer encoder along with their label encodings. The transformer encoder learns to extract contextual information from support images and predict the query image in a single pass, eliminating the need for the fine-tuning step during inference. An overview of our approach is visualized in Fig. 1.
@@
Figure 1: Visualization of CAMeLU (with 3-way 5-shot tasks). The left side illustrates the task creation mechanism, where $N$ samples are drawn from an unlabeled dataset $\mathcal{D}_{train}$. Each sample $x_{n}$ is augmented $K$ times to obtain $x_{n,k}^{(sp)}$. A strategy inspired by mixup (Zhang et al., 2018) is utilized for generating the query set by using an augmented version of $x_{n}$, i.e., $\tilde{x}_{n,j}$. The same pseudo-label $n\in[1,N]$ is assigned to all data generated from the sample $x_{n}$. On the right side, the so-created task is fed into the transformer encoder for predicting the query input. Inspired by CAML (Fifty et al., 2024), the transformer encoder processes demonstrations created by concatenating features from a fixed pre-trained feature extractor and a learned class encoder. The symbol * denotes the unknown query label that the transformer encoder aims to predict.
@@ 2 Related work
Meta-learning is a well-studied field in the machine learning community due to its ability to enable models to quickly adapt to tasks with limited labeled data. Pioneering work in the field (Finn et al., 2017; Snell et al., 2017; Vinyals et al., 2016; Mishra et al., 2018; Sung et al., 2018) considers the scenarios where a large labeled dataset is available for meta-training, a challenging requirement in real-world applications. UML addresses this challenge by extracting meaningful information from unsupervised data that can be transferred to downstream tasks with limited labeled data. Different techniques have been explored in the literature to construct diverse tasks. CACTUs (Hsu et al., 2018) applies clustering in the embedding space and assigns the same pseudo-label to all images in the same cluster. Other methods focus on generating synthetic samples, either using data augmentations, as in UMTRA (Khodadadeh et al., 2019), or leveraging interpolation in the latent space of a generative model (Khodadadeh et al., 2020). Differently, Meta-GMVAE (Lee et al., 2020) and Meta-SVEBM (Kong et al., 2021) use variational autoencoders and memory-based models for pseudo-label generation. Recent methodologies have also incorporated SSL techniques (Doersch et al., 2015) into UML methods. In particular, Set-SimCLR (Lee et al., 2022) builds on top of the SimCLR (Chen et al., 2020) approach and reframes meta-learning as a set-level problem, while PsCo (Jang et al., 2022), inspired by MoCo (He et al., 2020), utilizes a momentum encoder and a queue of previous samples to improve pseudo-labeling and construct diverse tasks for UML applications. Similarly, BECLR (Poulakakis-Daktylidis & Jamali-Rad, 2024) introduces an approach for unsupervised few-shot learning by proposing a constrastive representation learning framework, instead of meta-learning.
@@
Several UML approaches rely on data augmentation to construct the training tasks (Khodadadeh et al., 2019; Lee et al., 2022; Jang et al., 2022). However, traditional transformations such as rotation, translation, cropping, resizing, and flipping (Shorten & Khoshgoftaar, 2019) might generate images that are too similar to the original ones, ending up in tasks with low in-class variability between the support and query images. This creates a problem when the model needs to generalize to test tasks, where the query data are different instances than the support ones, not only augmented versions of them. In this paper, we addressed this limitation by generating query images using a strategy inspired by mixup (Zhang et al., 2018) to enhance model generalization. Similarly to mixup, which performs linear interpolation of the feature vectors at the pixel level, other strategies based on mixing images comprise CutMix (Yun et al., 2019), PatchMix (Liu et al., 2021), and Manifold Mixup (Verma et al., 2019).
@@
In-context learning refers to the ability to perform a new task via inference alone by conditioning on a few input-output pairs and making predictions for new inputs (Dong et al., 2022). Although typical of LLMs (Devlin et al., 2019; Radford et al., 2019; Touvron et al., 2023), this ability has also been explored in different fields, such as in-painting (Bar et al., 2022; Zhang et al., 2024), image segmentation (Butoi et al., 2023), and notably meta-learning (Chan et al., 2022; Singh et al., 2024; Kirsch et al., 2022; Fifty et al., 2023; 2024; Min et al., 2022). Recent methods, such as Chan et al. (2022) and Singh et al. (2024), examine the emergence of in-context learning abilities from a data distribution perspective, extending these insights to images. GPICL (Kirsch et al., 2022) further demonstrates that transformers can be meta-trained as general-purpose in-context learners, while CAML (Fifty et al., 2024) adapts this concept to non-causal sequence modeling problems. Building on these advancements, our work takes a different direction by tackling the unsupervised meta-learning problem. Specifically, we introduce a novel task creation mechanism that, together with an in-context learner, enables learning directly from an unlabeled dataset. This approach differentiates our method from prior in-context learning techniques, aligning it with the unique requirements of UML.
@@ 3 Proposed approach
Our proposed approach, Context-Aware Meta-Learning in Unsupervised scenarios (CAMeLU), leverages the in-context learning ability of transformers to address the challenges of UML. These challenges include the need to construct meaningful tasks from unlabeled data and the requirement for models to generalize effectively to new tasks during inference. CAMeLU consists of two phases that are intertwined during the model training. Initially, tasks are automatically constructed from an unlabeled dataset utilizing a combination of two strategies. Subsequently, we reformulate the meta-learning framework as a sequence modeling problem, aiming to harness the in-context learning capability of a transformer. This enables the model to extract context from the support samples and predict the unknown query samples without requiring any fine-tuning during the inference phase. The combination of these two phases is essential and guarantees good generalization performance without labeled information. Transformers excel at modeling dependencies and capturing relationships between support and query samples, which is particularly beneficial in few-shot learning scenarios. The novel task creation mechanism complements this by constructing diverse and challenging pseudo-tasks, effectively preparing the model for the complexities of target tasks. We delve into the two phases in Sect. 3.1 and Sect. 3.2, respectively.
@@ 3.1 Task creation
For the support set, we randomly sample $N$ images from $\mathcal{D}_{train}$ under the assumption that they belong to distinct categories, as shown in Fig. 1. This assumption is reasonable when $N<<C$, where $C$ denotes the total number of classes in $\mathcal{D}_{train}$, which is satisfied using a large training dataset. If we assume that all samples are equally distributed among the classes, i.e., $m$ samples per class, the probability that two or more samples are in the same class is equal to
@@
This task-creation mechanism can be seen as a task augmentation strategy (Yao et al., 2021; Rajendran et al., 2020) that allows the generation of a large number (almost infinite) of diverse tasks. This is particularly useful for meta-learning and in-context learning applications where the model needs to acquire knowledge from a multitude of tasks to generalize to unseen tasks sampled from different domains.
@@
While the strategy used for generating the query images draws inspiration from the mixup strategy proposed in Zhang et al. (2018), there are some substantial differences. The aim of mixup is to develop a new data augmentation strategy to expand the number of training examples and diversify the data distribution used for training, thereby enhancing the robustness and generalization of neural networks. In CAMeLU, the primary objective of merging images is to encourage the model to learn even in scenarios where only a fraction of the class context is present in the image. In CAMeLU, $\lambda$ is sampled from a uniform distribution (obtained with a $Beta$ distribution with $\alpha=1,\beta=1$) in $(0,0.5)$, guaranteeing that the amount of information from $z_{j}$ that is embedded into $x_{j}^{(qr)}$ is less than 50%, thus ensuring that the assigned label is consistent with the class of the support images generated from $x_{n}$. Indeed, we assign the same label $n$ to $x_{j}^{(qr)}$, forcing the network to learn to retrieve information in $x_{j}^{(qr)}$ that is related to the category of $x_{n}$. Contrarily, mixup creates new examples by interpolating both images and labels at the scope of limiting memorization over the training distribution.
@@ 3.2 In-context learning method
Following task creation, we rephrase the meta-learning framework as a non-causal sequence modeling problem, where the order of the examples does not entail a causal relationship. Inspired by recent developments in LLMs (Garg et al., 2022; Li et al., 2023; Devlin et al., 2019; Radford et al., 2019; Touvron et al., 2023), we treat each task as a prompt, where the support embeddings, together with the learned projected labels, form the demonstration context, whereas the query represents the classification problem that the network is required to solve. A model is said to in-context learn a task if it can approximate $y_{j}^{(qr)}$ for a new query input $x_{j}^{(qr)}$ by conditioning on a sequence $S_{i,j}$ containing in-context (support) examples and one query input defined as follows:
@@
To achieve this, we design a model comprising three components: (1) a feature extractor $f_{\psi}$, (2) a class encoder $g_{\phi}$, and (3) a transformer encoder with a linear projection layer on top, i.e., $M_{\theta}$. The feature extractor aims to map support and query samples into a latent space where images with similar characteristics and semantic meaning are assigned similar representations. In Appendix A.4 we explore various feature extractors for this purpose, including those pre-trained via a supervised approach or leveraging an SSL technique. The resulting representations are then concatenated with a class embedding. The class embeddings for the support representations are generated by encoding the corresponding classes using the class encoder $g_{\phi}$. However, as the classes of the queries are unknown, a randomly initialized learnable vector is appended to each query representation. The so-combined embeddings are then organized into sequences $S_{i,j}=\left((f_{\psi}(x_{1}^{(sp)}),g_{\phi}(y_{1}^{(sp)})),\ldots,(f_{\psi}(x_{NK}^{(sp)}),g_{\phi}(y_{NK}^{(sp)})),f_{\psi}(x_{j}^{(qr)})\right),$ $j=1,\ldots,Q,$ resembling the one in Eq. 2. These sequences are fed into the transformer encoder, and only the transformer output corresponding to the query sample is selected and passed through a projection layer to predict the query label. This process iterates for all queries in the task, and the aggregated loss is employed for model training. In particular, the training process can be formulated as an optimization process where the objective is as follows:
@@
During evaluation, when a new task is presented, the available examples in $D_{new}^{(sp)}$ are utilized as contextual information to guide the classification of the query samples without requiring any fine-tuning or adaptation steps.
Analysis.
(a) Feature extractor
(b) Transformer encoder
Figure 2: Visualization of clustered embeddings obtained with CAMeLU after the feature extractor (left) and the transformer encoder (right) on a 5-way 5-shot task sampled from the CUB dataset. Crosses indicate the centroids of each class, and the numbers denote the Euclidean distances between the query (triangle) and each class centroid. The plots are obtained using t-SNE (Van der Maaten & Hinton, 2008) with a perplexity equal to 9.
To gain a better understanding of how the in-context learner functions during inference, we show the embedding space of an exemplary test task after the feature extractor and the transformer encoder in Fig. 2. The embedding space after the feature extractor appears sparse, with a large Euclidean distance between the query sample and the centroid of each class, which indicates limited class separability and less informative representations. In contrast, the transformer encoder significantly improves the representation, producing more compact and well-separated clusters. Notably, the query representation aligns closely with the support examples of the same class, showcasing the effectiveness of the transformer in utilizing the task context to refine predictions. This demonstrates the in-context learner’s ability to adapt representations dynamically based on the task context and also provides evidence supporting CAMeLU’s superior performance, particularly in cross-domain and few-shot scenarios where generalization is more challenging. Experiments for the other datasets can be found in Appendix A.2.